### v1

# DLPOLY 5: Calculation of system properties on the fly for very large systems via massive parallelism

Abstract
Modelling has become a third distinct line of scientific enquiry, alongside experiments and theory. Molecular dynamics (MD) simulations serve to interpret, predict and guide experiments and to test and develop theories. A major limiting factor of MD simulations is system size and in particular the difficulty in handling, storing and processing trajectories of very large systems. This limitation has become significant as the need to simulate large system sizes of the order of billions of atoms and beyond has been steadily growing. Examples include interface phenomena, composite materials, biomaterials, melting, nucleation, atomic transport, adhesion, radiation damage and fracture. More generally, accessing new length and energy scales often brings qualitatively new science, but this has currently reached a bottleneck in MD simulations due to the traditional methods of storing and post-processing trajectory files. To address this challenge, we propose a new paradigm of running MD simulations: instead of storing and post-processing trajectory files, we calculate key system properties on-the-fly. Here, we discuss the implementation of this idea and on-the-fly calculation of key system properties in the general-purpose MD code, DLPOLY. We discuss code development, new capabilities and the calculation of these properties, including correlation functions, viscosity, thermal conductivity and elastic constants. We give examples of these on-the-fly calculations in very large systems. Our developments offer a new way to run MD simulations of large systems efficiently in the future.
keywords:
, Molecular dynamics , Materials modelling , online algorithms , DL_POLYPROGRAM SUMMARY/NEW VERSION PROGRAM SUMMARY
Program Title: DLPOLY_5
Developer’s repository link: https://gitlab.com/ccp5/dl-poly
Licensing provisions: L-GPL v3.0
Programming language: Fortran95
Supplementary material:
Journal reference of previous version: Todorov, I.T., Smith, W., Trachenko, K. and Dove, M.T., 2006. DL_POLY_3: new dimensions in molecular dynamics simulations via massive parallelism. Journal of Materials Chemistry, 16(20), pp.1911-1918.
Does the new version supersede the previous version?: Yes
Reasons for the new version: Development of on-the-fly correlation analysis.
Summary of revisions: General code modularisation, a new control format, SPME re-implementation for per-particle contributions, and a general on-the-fly correlation framework.
Nature of problem: Molecular dynamics is utilised for modelling in many domains including physics, chemistry, biology, materials science, and their interfaces. These applications often call for large scale simulations targeting high-fidelity timescales and ever larger length scales. In all these cases efficient use of computation, storage, and input/output (I/O) load handling on both the software and hardware sides is required to facilitate analysis.
Solution method:
DL_POLY provides an efficient set of algorithms for molecular simulation alongside a domain-decomposition strategy to distributed computation efficiently across massively parallel CPU systems. These include parallel I/O handling, link-cells, smooth particle mesh Ewald electrostatics, and now a general purpose on-the-fly correlation framework. This latter development addresses the problem of storing and analysing large trajectory files by iteratively computing correlations at runtime.
1 Introduction
For a few decades now, molecular simulations have grown into a third distinct line of scientific enquiry in condensed matter, alongside experiment and theory [1, 2]. Molecular dynamics (MD) simulations give coordinates and momenta of simulated particles as a function of time and thus provide the system’s trajectory in phase space. This enables us to calculate most important system properties including those related to structure, statistics, dynamics, transport and so on.
MD results are used in a number of ways to provide insight into experiments, from interpretation to prediction and guiding new ones, as well as to test theories and provide insight at the atomic level resolution for materials, scenarios or processes not yet synthesised by researchers or constructed by experimentalists. MD is particularly useful in fields where measurements are inaccessible in experiment, where empirical theory of phenomena at the larger engineering scales can be examined at short time and length scales, such as damage in materials [3]. Another popular use is extracting statistics and analytics that are impossible to capture by experiment such as chem-informatics analysis (formation of hydrogen bond networks, hydrophobicity, - stacking, etc.), the dynamics of defects in materials (defects and faults in crystalline matrices, amorphisation processes) as well as coarse-grained modelling using augmented force-fields of complex dynamics of, for example, proteins. Generally, MD results are also used to generate synthetic benchmark data for fortifying machine-learnt (ML) models.
This interaction between experiments, theory and simulation has proven to be very productive and has naturally found wide applications in physics, chemistry, materials and earth sciences, engineering, biological and medical sciences etc, from purely academic research to real-world industrial applications. These applications include those in the strategic priority areas of energy, environment, advanced materials and health. It is hard to imagine the current state of these areas without MD and other types of computer modelling. Many practising scientists have taken MD simulations on board to the extent that they consider themselves as combined experimentalist-modellers or theorist-modellers.
As for any method, MD simulations have their limitations and associated domain of applicability. Several important ones have been relaxed or lifted through the developments of the last decades; others remain unsolved, including the important issue of small system size.
One of the strengths of MD is its ability to access small length scales of nanometers and short time scales of picoseconds or beyond where few experimental methods can measure. Early MD simulations were able to simulate about 100 particles for typically tens of picoseconds. With increasing computer power, simulations of millions of atoms on a personal computer is now routine, with simulation times extending to nanoseconds and above.
The appetite of the modelling community to simulate larger systems is constantly growing. In many areas, large systems are necessary to reach the required length or energy scales and match these to experiments and relevant physical processes. These areas include interface phenomena, composite materials, biomaterials, melting, nucleation, atomic transport, adhesion, radiation damage and fracture.
About a decade ago, we simulated system sizes approaching 1 billion atoms with realistic many-body potentials for iron, using over 60,000 parallel processors [4]. This large system size was required in order to contain radiation damage due to high-energy (in MeV) recoil energy. A representative picture of the damage is shown in Fig. 1. We note that the system size, in the order of micrometers, is in interesting proximity to the length scales visible through an optical microscope; simulations and experiments met at this length-scale for the first time.
We found that the nature of the damage and defects in these structures were new and could not be extrapolated from lower-energy events. This had a clear value from scientific perspective and showed that, as is often the case in physics, exploring new length- and energy-scales brings about new effects [4].
Running these very large simulations, not only brought scientific insights, but also computational insights. First, the size of a single configuration of 100 million atoms (position data only, in Å with one significant figure) in ASCII is 3.5GB. For comparison, the standard DLPOLY format requires 5 times more data for structural data (15 GB) with an additional 2 times more if velocity data is needed (25GB). In an MD simulation, certain structural and dynamical properties (for example, viscosity or velocity correlation functions) often need over 10,000-100,000 configurations to be computed, requiring usually two orders of magnitude more timesteps in order to get good-quality data. This means that a single trajectory file of this type for a 1 million atom system will require 0.1-1 PB of storage. On the other hand, the work file system of a high-performance facility such as UK’s ARCHER2 service [6] has a total of 10.2 PB [7] to be shared between all users (a maximum of 60 TB allocated to a single project or team). Therefore, the trajectory files of very large systems become impossible to store, let alone process, on high-performance computing facilities.
Second, a high-performance computing facility such as ARCHER2 spends a significant amount of its time writing the data rather than actually carrying out the numerical simulation, which is an inefficient way of utilising the HPC resource. Even with efficient parallelisation of writing data to storage, a best outcome means that writing a single sample configuration costs as much in CPU time as a timestep. In the case of 100 million atoms run on 65,000 processors, this fraction was in excess of 5% for a full production run, including reading and writing. This implies that each simulation run involving writing the trajectory file, even if feasible in terms of file size, wastes considerable compute power. Furthermore, the larger the simulation (in model size and time evolution) and the larger the count of processors used to run it, the larger the waste.
The problems outlined above are general and will either prevent users from running and analysing MD simulations of large systems or make these simulations inefficient. In our radiation damage case study, we overcame these problems by implementing a special on-the-fly algorithm in the general-purpose MD package, DLPOLY [8]. This algorithm specifically finds and analyses displaced and defect atoms in the very large billion-atom system. This was a non-trivial task as the algorithm had to be made consistent with the domain-decomposition infrastructure of DLPOLY [9, 10]. These simulations have provided a useful case study and a motivation for the development of a more general code, with a view to benefit a wider modelling community and to approach the problems in a general way.
We recall that some properties such as energy can be sampled and averaged at each time step, and these are calculated by DLPOLY and other popular MD codes as a standard output. However, there are other important properties relying on time correlation functions which need to be calculated over the entire trajectory and which are currently calculated by writing out the trajectory file and analysing it a posteriori. One common example is the velocity autocorrelation function related to the phonon density of states and power spectrum of the system. Other examples include k-space densities and currents. For very large system sizes, the problems discussed above make performing these analyses from trajectory files impractical.
Instead of attempting to run a trajectory and subsequently analysing it, we have enabled the calculation of important physical properties on-the-fly.
This sort of on-the-fly calculation will be crucial to the execution and analysis of ever-larger MD simulations. Particularly as there is a growing need to simulate larger systems, length and energy scales, coupled with the availability of scalable MD codes and massive parallel computing facilities.
Supported by the EPSRC funding, we have extended the UK flagship MD code, DLPOLY, to calculate system properties on-the-fly culminating in the DLPOLY 5 code available at [11]. DLPOLY is a general-purpose massively parallel MD simulation package that uses a highly efficient set of methods and algorithms, including domain decomposition, linked cells, Daresbury Advanced Fourier Transform, Trotter derived Velocity Verlet integration and RATTLE [8]. Written to support academic and industrial research, DLPOLY has a wide range of applications and can run on a wide range of computers; from single processor workstations to massively parallel high-performance facilities. During development a strong emphasis was placed on efficient utilization of multi-processor power by optimising memory workload and distribution. This makes it possible to efficiently simulate large systems of billions of atoms. Simulation of systems of this size has been tested, and we are not aware of reasons preventing from running larger systems. DLPOLY 5 is hosted on GitLab [11] has been a free and open source project (GPLv3.0) since 2020.
In this paper, we document and detail the developments of DLPOLY related to on-the-fly calculation of some key system properties. This included general and substantial refactoring across DLPOLY’s modules. We discuss the implementation of on-the-fly calculations of different properties including correlation functions, viscosity and thermal conductivity, elastic constants, currents, and rigid body properties. We then discuss the calculations of these properties and show the results from simulations of large 100-million systems. These results are compared to smaller-size calculations, experiments and previous MD simulations. We begin with an overview of DLPOLY’s history up to DLPOLY 4, including a discussion on its scalability. Then we detail the re-engineering of the code, laying the groundwork for our on-the-fly developments.
2 DLPOLY, massive parallelism and scalability
The DLPOLY project was conceived in 1993 by W Smith and originally released in 1994 as DLPOLY 2 [12, 13] (now known as DLPOLY_CLASSIC [14]) using a replicated data (RD) approach as its underlying parallelisation strategy [15, 16]. Written in Fortran77 [17] and using MPI [18] for interprocessor communication the project brought significant improvements in both functionality and internal algorithms performance in the code, establishing the link-cell (LC) algorithm for efficient search and listing of particle pairs based on a cutoff [19] and the use of 3D discrete fast Fourier transforms (3D FFTs) [20] for the central grid-based operation in the evaluation of the long-ranged contributions of electrostatics interactions [21]. However, by the early 2000s it was clear that RD strategy was rather limited in performance and scalability gains especially for then very large model systems, of the order of 100,000 atoms, and large processor counts, of the order of 100 CPUs, thus necessitating a change in parallelisation strategy and reformulation of its algorithms. The particular deficiencies of the RD approach that needed addressing were the avalanche of communication which increased with the size of a model system (creating larger inter-processor messages) as well as with the processor count (increasing the number of senders and receivers per message for all-to-all communications), and that there was no memory distribution at all, these were major bottlenecks for running large systems of the order of millions of atoms.
In 2003, led by I. Todorov, with core work carried out by W. Smith and I. Bush, DLPOLY 3 was released as part of the fundamental work supported by the eMinerals project [22], brought in as part of the UK’s eScience programme [23]. Not only did this version bring a new code infrastructure in modularised Fortran 90, but it also developed the foundation of the domain decomposition (DD) parallelisation strategy [9, 10], including a Smooth Particle Mesh Ewald (SPME) method suitable for the DLPOLY DD [24], still used in the newer versions. Based on the two premises of: (i) limited density variation - that the model system average particle density does not fluctuate significantly - in space and time, and (ii) the system size/dimensions being much larger than that of the most characteristic interaction cutoff; this strategy developed and addressed the two major shortcomings of the RD approach. First, it provided for near perfect memory distribution as demonstrated in Fig. 2, where three model systems with increasing complexity were run and compared in a weak-scaling manner, i.e. every time the system size doubled, the processor count used to run the simulation doubled. This opened the door to running multi-million atom simulations from its very first release [25].
The weak scaling provides information about the deviation from perfect performance, where it is assumed that doubling processors halves the time-per-atom, i.e. constant speed gain. However, we see in Fig. 2, that even for the most trivial system, Ar, one can see small deviation from the perfect performance due to inter-process communication, necessary for domain halo boundary data exchange. It is clear that, as the model’s complexity increases, so too does the toll on performance, demonstrated by the greater deviation from perfect scaling. For the second model system, NaCl, this is due to long-ranged coulombic interactions, bringing the cost of inclusion of electrostatics into focus. As discussed [24] this is due to communications overheads that are a necessity of the FFT operations needed for evaluation of the long-ranged electrostatics interactions. The third model system, SPC water, is subject to a further performance penalty which is caused by the inclusion of constrained bond dynamics. The reason for this is simple to understand: as more processors are working together, there are more inter-domain faces and hence more constrained oxygen-hydrogen bonds crossing inter-domain faces. Constraining these bonds requires multiple small inter-process updates, which become costly as the model system or processor counts become bigger.
In strong scaling, we track the performance of the key underlying algorithms on the same system size as we increase the processor count. As shown in Fig. 3 the domain decomposition parallelisation strategy provides a good load-balancing and near-perfect (linear) performance, up to to a limit, by the majority of its core algorithms. The speed gain curve deviates from its linear form only for the long-range interactions (Ewald k-space). This evaluation is also the ultimate cause for the cumulative overall speed gain to deviate in a non-linear manner as the processor count increases and decay with increasing the processor count. More recent strong-scaling benchmarking on Archer 2 [6] demonstrated the same excellent scalability for large Ar and NaCl systems. A comprehensive survey of DLPOLY performance on modern hardware can be found in [26].
In 2010 DLPOLY_4 was released, including a major rewrite providing an extension of particle dynamics to rigid body dynamics [27], inclusion of two-temperature thermostat functionality [28] for high-energy events and the inclusion of parallel I/O [29]. The rigid body dynamics provided a natural extension of the parallel performance capabilities for biological and biochemical systems in which constrained bond dynamics was a bottleneck (as seen in the fast decaying weak performance of the SPC water system in Fig. 2). The enhancement in performance of rigid body (RB) dynamics and constrained bond (CB) dynamics is demonstrated in Fig. 4 where both weak and strong scaling tests were performed on the same SPC water system and model with scaled sizes in the case of weak scaling. The difference between RB and CB dynamics is the way in which bond constraints are applied. In the CB case, it is based on simple, but iterative, application of small constant force updates on top of the Newtonian equations of motion, which in DD requires frequent local inter-process data exchanges. In contrast, in the RB case the Eulerian equations of RB rotation are solved alongside with the Newtonian ones. However, these further calculations, despite their complexity, are computationally cheaper (no iterative computation) and require no iterative communication for a difference with CB dynamics one. Further developments in 2012 led to the inclusion of Dissipative Particle Dynamics integration [30] and a three seeded random number generation [31] for using repeatable random series for stochastically dependent thermostating, kinetic fields and pressure tensors in a fully deterministic manner across domains and time series.
2.1 Current state
A few years ago an extensive refactoring effort took place that is the the base of current DLPOLY. These changes were introduced mainly due to the need for being able to add new complex tasks with maximum code reuse and allow various easy optimisations. The main changes introduced were: i) using object oriented Fortran – 2018 as the minimal standard required, ii) restructuring the code to be fully modular, with all functions and subroutines exchanging data only via parameters rather than via modules - removal of implicit common blocks. Only one singleton remains for dealing with I/O for errors. iii) introduced an automated test system with both regression and unit tests. Testing was extended from 24 tests to over 200 tests. iv) atomic configurations are stored in an array of structures that is cache friendly with respect to positions and forces, v) electrostatics was rewritten to allow per-particle decomposition and allowing the computation of long range contributions from any power potential, see next section for details. iv) The CONTROL file format was redesigned (see appendix C). In general all entries now are consistently of the form token-value-unit (where applicable), STATIS, RDF and other output files can be written in the yaml format [32] for easier postprocessing, v) a modern software infrastructure is in place with full CI/CD via gitlab.com, vi) license was changed to an open-source license GPL-3.0, vii) dlpoly-py was developed [33] - a companion python package that can read inputs and outputs of DLPOLY and initiate simulations. Without this extensive refactoring, implementation of the Empirical Valence Bond [34] and the new Chemshell interface [35] would not have been possible.
2.2 SPME refactoring
The Ewald method [36] is a method for the calculation of long-range (conditionally convergent until infinity) forces. This can be achieved in a number of ways, the usual physical description being the introduction of a screening charge distribution and a correction to account for the screening so as to accelerate the summation of the long-range component in Fourier space. This can also be seen as performing a sum from in real-space and from in reciprocal space to construct a full sum of in finite time ().
The smooth particle mesh Ewald (SPME) approach by Essmann et al. [21] uses cardinal B-Splines to interpolate the complex exponentials that result from the above approach in a regular fashion. This allows the use of a standard Fast Fourier Transform (FFT), so reducing the scaling of the method with system size to . However this is at the cost of some accuracy due to the interpolation, though this can be negated by using higher order splines to provide a better description of the exponentials.
The SPME approach in DLPOLY has been refactored to enable computation of per-particle contributions to energies, forces and stresses. The calculation of per-particle contributions has been used to implement the computation of Green-Kubo thermal conductivity with Ewald methods (See section 3.3).
This new method coincidentally enables the expansion to not only alternative orders of potentials, but also derivatives of arbitrary order, thereby paving the way to correct multipolar expansions of arbitrary order. Direct computations of higher-order potentials allow “exact” computation of, e.g. Lennard-Jones (6-12) [37] potentials without cut-offs, enabling calculation of the effect of these cutoffs on actual dynamics. Currently implemented are functional forms for: , , , and a general form for and their derivatives for both the real- and reciprocal-space parts (see appendix A).
As part of the DLPOLY5 developments, we have recently benchmarked 2-body and SPME-enabled systems composed of million atoms and 2-body simulations composed of billion atoms across increasing MPI process counts on Archer2 [6]. We examine processor counts ranging from 128 in the small systems to 262,144 in the larger system. Representing a 35% concurrent utilisation of Archer2’s 750,080 CPU cores [38]. The results can be found in Fig. 5. Primarily the results continue to demonstrate the scalability of DLPOLY. But secondly, it also demonstrated the scalability of the newly refactored MD algorithms, with the reworked DD memory distribution that favoured purely local communications and minimised all-to-all communications as much as possible. A comprehensive survey of DLPOLY performance on modern hardware can be found in [26].
3 System properties on-the-fly
3.1 On-the-fly methods
Transforming data as it is acquired is a paradigm often termed “online algorithms” [39]. The basic idea is that, instead of processing an entire data set a once, an on-the-fly calculation processes the data incrementally, often as the data are generated, received, or simply by processing a data set in chunks. This can incur performance advantages by eliminating the requirement to store the full data set on disc or in memory (RAM). Alternatively input/output (I/O) penalties can be significant when reading from disk, and RAM may be limited. In some cases, the full data set may be either too large to store in memory, or perhaps even unbounded. One example of unbounded data is the internet, where a continuous stream of “big”-data is encountered on a daily basis [40]. In these applications a sliding-window technique is often used where statistics and other data transformations are defined in reference to a fixed-size window [41] which could be temporal or spatial.
In MD, the configuration of atomic positions is mapped through simulation time-steps into a trajectory. In order to calculate any properties defined by these time-linked configurations the trajectory may be stored in memory, and an algorithm applied. For example storing atom velocities and calculating the velocity auto-correlation function (VAF). For small systems, it may be possible to store these data at every step, obtaining a VAF accurate to the time-step . As the system size scales this becomes infeasible both in terms of storage and I/O penalties if storing to disc. By conceptualising the simulation trajectory as a big-data stream, we can apply the same ideas of online algorithms to calculate system properties on-the-fly. For modest systems, this presents a convenience. However for large systems, this methodology can give access to properties which otherwise require infeasible storage requirements or some accuracy trade-off.
Multiple system properties depend upon the calculation of correlation functions. As an incomplete list: (1) The VAF is one example which can in turn be used to calculate the vibrational density of states. (2) From Green-Kubo theory [42, 43] the thermal conductivity and viscosity can be calculated by using correlation functions of the heat flux and stress tensor respectively [42, 43]. (3) Elastic constants may also be calculated using either strain correlations [44] or stress correlations [45, 46]. (4) k-space density correlations may be used to determine thermal-conductivity [47], and momentum current correlations may be used to calculate structure factors and liquid phonon spectra [48, 49].
Motivated by these examples, and others, in DLPOLY we implemented a general on-the-fly correlation algorithm, so that we are able to calculate these values at runtime; that is as the system is simulated, without storing trajectory data to disc.
3.2 Correlation functions at runtime
Given two observable quantities and in MD, where are simulation time steps for a time step of , we can evaluate the correlation function of these two values at time-lags as [43],
for one trajectory. Typically statistics are accumulated over multiple trajectories. In any case, directly evaluating equation 1 is cumbersome as and increase. Even more so if the observable is averaged over atoms of molecules (for example, in the case of the VAF). Therefore, often the discrete Fourier transforms of and are taken to access the correlation function using the inverse transform of their product [43]. Whether the sum is taken directly or a Fourier transform method is used and must be stored in memory or on disc for the computation.
We use the multiple-tau correlation algorithm [50] within DLPOLY to calculate correlations on-the-fly. The method works by maintaining an hierarchical set of blocks which are updated during the simulation. The blocks store observed data (velocities, stresses, etc.) in a decreasing resolution. The first block stores the observed data intact whilst lower blocks store block averages, computed over the previous level. This is controlled by three parameters, the number of blocks , the number of points within each block , and the size of the block average . See Appendix D for algorithmic details of our implementation based upon [50].
This approach is similar to the velocity correlator introduced by Frenkel and Smit [2] with the additional separation of (data points stored at each level) and the averaging parameter. By increasing and (with fixed), later correlation times may be collected at a fixed accuracy. Increasing has the effect of calculating much longer correlation times for a trade off in accuracy and increased performance. The maximum lag-time of a correlation given these parameters is defined by where is the simulation time step and the frequency the correlation is updated is . The trade off in performance terms scales as . The actual performance will also depend on the values correlated (e.g. velocity across all atoms, or system stress) and other system options (such as electrostatics).
In DLPOLY we support correlating key quantities such as components of heat-flux, velocity, the stress tensor, rigid body positions, velocites, and angular velocities, as well as -space resolved densities, heat-fluxes, currents, stress tensors, and energy currents. Additionally common statistical values reported by DLPOLY’s statistics module including, but not limited to volume, kinetic/potential energy, and temperature. We do this by defining each as an observable. Quantities that are observable may be juxtaposed in the CONTROL file to define arbitrary correlations. For example a velocity-velocity correlation may be defined by v_x-v_x where _x indicates the component correlated, and stress_xy-stress_xy defines a stress correlation of the -component of the stress tensor. Scalar quantities such as volume can be requested using e.g. volume-volume. The power of the juxtaposition is that any pair of observables can be correlated.
For each correlation, independently, we also support specifying the parameters , , and separately, and provide the ability to set the frequency each correlation is updated separately from the frequency of other statistics and output in DLPOLY. This enables high-resolution correlations to be calculated without additional I/O overhead, and without forcing low-resolution correlations to be calculated at unnecessarily high time-resolutions. Differences in correlation frequency are handled in the correlation output by automatically accounting for the different effective time-units in the outputted lag times.
During the simulation, correlators are stored index-able by the unique names of the observables they correlate, i.e. at any point in the simulation, we may trivially obtain the current correlation value for say the heat-flux in , heat_flux_x-heat_flux_x. Not only does this serve to maintain uniqueness, but also allows us to easily report these values, or process them further, at runtime. The result of each correlator is reported in a new file COR which reports the value of each correlation function as well as potentially a set of derived quantities, at user specified intervals. These derived quantities include viscosity, thermal conductivity, and elastic constants as detailed in the next sections.
3.3 Viscosity and thermal conductivity
The viscosity and thermal conductivity are important properties characterising liquids, for example in the evaluation of the performance of molten salts as nuclear reactor coolant fluids, since this is determined by their transport and thermal diffusion characteristics [51, 52]; these can be calculated using Green-Kubo theory from correlations of shear-stress and heat-flux respectively [42, 43, 53].
The shear-viscosity may be calculated using the integral
where is the ensemble average, is the system temperature and is Boltzmann’s constant. Finally the stress tensor, for pairwise additive potential , is defined as
where , , are particle ’s mass, velocity, and position, and , and is the system volume. Finally, and are orthogonal Cartesian coordinates.
Similarly the thermal conductivity is related to the heat flux by
Where the heat flux is
where is the total energy (kinetic plus potential) of particle , and is the force exerted on due to . In Eq. 4 the dot product has the function of averaging over the dimensions , , and .
For multi-component systems with significant asymmetry (e.g. mass difference) it is necessary to correct for mass flux effects when calculating in order to compare to experimental data [54]. For example the effect is significant in LiF but not in KCl [55] which have mass ratios and respectively. This can be done by forming multiple Green-Kubo formulae using the heat flux and partial momentum densities for particular species,
where is the set of atom indices for atoms of species . In DLPOLY the partial momentum densities may be calculated, and correlated, for a selection of species by the user. The values themselves may also be written to the HEATFLUX file alongside . In general it is possible to take advantage of momentum conservation in order to not calculate all partial momentum densities. A user is then able to form the necessary correlations for the corrected Green-Kubo formula required.
3.3.1 Comparison to experiment
The correlation functions integrated in equations 2 and 4 can be calculated during an MD run using DLPOLY’s on-the-fly correlator. In each case we can approximate the integral by computing the required correlation to some maximum time-lag and performing a numerical integration of the correlation functions’ values. In each case, when multiple congruent correlation are present (e.g. and shear stresses), DLPOLY automatically calculates the derived properties (viscosity) listing them for each correlation and also as an averaged quantity. Care must be taken to collect adequate statistics for accurate calculations of these derived quantities. For example Zhang et al. [53] detail a method for choosing an optimal cutoff lag-time.
For argon, there is experimental data for and at a variety of pressure and temperatures compiled by the National Institute of Standards and Technology (NIST) [56]. DLPOLY has been used successfully before to calculated viscosity for argon by directly correlating the outputted stress tensor [57]. Here we report new data including thermal conductivity using DLPOLY’s on-the-fly correlator.
For all simulation work, we began by equilibrating for steps in an NPT ensemble, followed by steps with an NVE ensemble to account for changing thermostats, and then production runs, where statistics are calculated, ran for time steps. In all cases the time step was ps. For correlations we collect data for lag times up to ps. For the small systems () we collect data on independently seeded simulations for replicates to ensure adequate averaging of the Green-Kubo results. For the large-scale systems, we were able to simulate replicates for each data point.
Figure 6 shows the small- and large-scale system results for viscosity (see Eq. 2). We find good agreement with the NIST data as expected [57]. Additionally we find similarly consistent results for the large-scale simulations. Likewise for thermal conductivity, Eq. 4, we also find the calculated results are consistent with the experimental data (see Figure 7). The thermal conductivity data notably deviates more, this is due to increased noise between different replicates. For viscosity, we find standard deviations, across the different simulation results, of the order Pa for both system sizes, whereas for thermal conductivity the smaller systems’ standard deviations are of the order but for the large-scale systems this rose to at K and at K and K.
Both viscosity and thermal conductivity show an interesting feature: the minima which bound these properties from below. The minima are due to the dynamical crossover of particle dynamics [57]. Interestingly, the values at the minima are fixed by fundamental physical constants such as the Planck constant and electron mass [58, 59, 60].
3.4 Elastic constants
Another use case for the on-the-fly correlator is the calculation of elastic constants. In general, the elastic constants of a system are defined by the elements of the elasticity tensor . The values of which can be calculated, in full, using either the stress-fluctuation of strain-fluctuation methods, or by examining system stress after applying a set of strains. The stress-fluctuation method has been found to be more reliable and faster converging [61, 62, 63, 64].
For the calculation of elastic constants, we use this method where the elastic constants are given by,
where the parameters is the atom count, is the Kronecker delta, and is the microscopic stress tensor as in equation 3. Finally the Born term is defined as,
The components of the elasticity tensor can be reduced to independent values for physical media, and further to for physical media with cubic symmetry. Using Voigt notation we identify , , and . DLPOLY automatically reports the values of the elastic constants the 21 possible elastic constants when the required stress correlation terms are present. To compare with experimental data, from the elastic constants we can derive the bulk and shear modulus. The bulk modulus is
For the shear modulus there is a choice of averaging, (Voigt) and (Reuss) which give upper and lower bounds of the experimental results [65]. These take the form
Both and are reported for experimentally observed argon crystals. Note that the isothermal compressibility is often reported where , and the bulk modulus itself may also be more simply calculated from fluctuations in volume in an NPT ensemble as [1]
3.4.1 Comparison to simulations and experiments
The following methodology was used in all our simulation results. An FCC Argon crystal was initialised and equilibrated in NPT, for steps (timestep fs) with a Nosé-hoover thermostat with thermo- and baro-stat couplings of ps and ps respectively, and a cutoff of Å. The potential was the Lennard-Jones potential with the same parameters as [61].
After equilibration production runs (NVT) were simulated for steps unless otherwise specified. For these runs, the atoms where re-scaled to the average simulation cell from the equilibration phase. Statistical averages were calculated on a rolling window of ps updated every step, and correlations computed for a maximum lag time of ps, also updated every step using DLPOLY’s on-the-fly correlators. Everything else was kept the same. For each replicate , this procedure was carried out with a different random seed for equilibration and production. The methodology is similar to [61], except they generate initial inputs through Monte-Carlo (MC) methods.
Fig. 8 compares DLPOLY’s calculated elastic constants with existing simulation results on argon FCC crystals, including MD and MC results, which are notably quite sparsely reported along the temperature axis. Previous simulation results are obtained from MC [45] and MD [62, 63, 61, 64]. We find that the values reported by DLPOLY are consistent with previous simulation results. The results from Pereverzev [63] and Squire et al. [45] follow broadly the same trend in temperature, and the remaining results reported at K are also consistent with DLPOLY’s.
Given that we find consistent results for the individual elastic constants against simulated data we now compare with experimental values. For the bulk modulus () these follow the same averaging scheme. DLPOLY’s results, Fig 9, bound the experimental values of Dobbs and Jones [66] as expected. The bulk modulus is also in good agreement with Peterson et al. [67]. Although there is a notable plateau trend at K, which is not captured in simulations. Not that Pereverzev [63] also finds a linear relationship at low temperature consistent with our results (see Fig 8).
3.5 Currents
In DLPOLY 5’s currents module is capable of calculating various collective properties resolved to user supplied -space vectors. To begin, the most basic functionality is the computation of -space density, as well as transverse and longitudinal currents. These are defined as [48]
Equation 13 may be used to determine the intermediate scattering function using the correlations
which can be used to determine thermal-conductivity in fluids [47], as well as the static structure factor as the correlation value. The dynamic structure factor may also be obtained using the Fourier transform of . Analogous correlations of and, may be used to determine transverse and longitudinal propagating modes. These collective properties relate to the overall motion of the particles of the system. Collective excitations form an integral and well known part of the theory of solid and gaseous states, but for liquids, which combine strong interactions with dynamical disorder, this has historically not been the case.
Alongside these currents, DLPOLY can also calculate the -space energy density, energy currents, and -dependent stress tensor.
The energy density is given by
and the -dependent stress and energy currents from equations 18 and 19.
A user may request correlations of combinations of these various currents (component wise). Each -point specified by the user will result in one correlation matching these requests in the CONTROL file. One use case for current correlations is calculating phonon dispersion curves. These dispersions can be calculated in DLPOLY via the calculation of current correlations
From these correlation functions, the spectra are found from the Fourier transforms and . The frequencies corresponding to maxima of these Fourier transforms of each wave-vector are interpreted as the spectrum of collective modes [48, 49].
In liquids in particular, it is possible to observe an interesting feature of spectra not seen in solids: -gap in the solid-like transverse spectrum meaning a threshold value below which no such propagating modes exist [68, 49]. The -gap can be observed by plotting the relationship between these maxima and .
To calculate these data in DLPOLY we first equilibrate an argon system of 4,000 atoms in NPT to the desired temperature. We then expand to other desired system sizes using DLPOLY’s nfold capability. We then calculate the correlations functions on-the-fly over simulation steps ( fs per step), for all physically reasonable -points, . Our correlations are resolved for points per block, at a correlation frequency of every steps (lags up to fs). The resulting correlation functions are first averaged across independent trajectories. The frequency maxima are then found by taking Fourier transforms of these averaged correlations. The result of this is shown in figure 10 where we see the increasing -gap for the transverse mode for temperatures of K and K.
3.6 Rigid body correlations
In DLPOLY rigid bodies can be defined as part of user input in the FIELD file. Molecules may have single or multiple rigid body components. During a simulation each rigid body has position, velocity and angular velocity data which may be correlated in the same manner as per-atom correlations. That is DLPOLY will determine distinct rigid body types based on distinct rigid body components within distinct molecular types, calculate requested correlations at runtime, and average these results for each.
Using this on-the-fly correlation functionality we are able to reproduce the results of Brodka and Zerda [69] for SF6. We do this by simulating rigid bodies of SF6 molecules with the same Lennard-Jones potential parameters. For equilibration we set the temperature to 296 K, and pressure to the experimental values listed for the pressure by Brodka and Zerda. We simulate in NPT for ps to ensure equilibration. The resulting densities were similar to the and g/cm3 listed in their results. We then rescaled our resulting configurations to match those densities exactly, and then proceed in NVE collecting statistics over ns simulation steps. Brodka and Zerda use equilibration steps and production steps at a higher time step ( fs vs. fs). Our results for the VAF are shown in Figure 11 which match well.
The VAF can also be used as a definition of the Frenkel line in supercritical fluids [70, 71], which marks the crossover between liquid-like and gas-like particle dynamics and thermodynamics. Using the rigid body VAF functionality we perform this analysis for rigid body CH4 molecules. We follow the process of Yang et al. [72], using the same potential [73]. We first equilibrate our system in NPT at the desired ( bar) and conditions for ps and then collect statistics over a further ps. As before we find the equilibrated simulation densities to be approximately the same as the NIST experimental data for supercritical CH4 [56], and for increasing there is a loss of the VAF minimum as shown in figure 12.
4 Conclusions
We have used the paradigm of on-the-fly calculations to develop a general purpose correlation module in DLPOLY 5. This functionality exposes key properties derived from simulation data which may be composed into correlation functions by the user. This method provides a convenient way to perform correlation analysis as well as eliminating the cumbersome I/O operations inherent in saving trajectory data for post-processing. It also presents the continuous availability of correlation function during simulation time. This saving becomes significant for high temporal resolution correlations defined at the atomic or molecular level such as the VAF.
DLPOLY continues to show the same scalability for large CPU counts as past versions. Our correlation implementation naturally takes advantage of the DD algorithm and multi-processing architecture used in DLPOLY to distribute the calculation of atomistic correlations. We aim to continue the development of on-the-fly methods in and beyond DLPOLY 5. For example, to date we have exposed rigid body molecular positions, velocities, and angular velocities. A natural extension is to correlate properties of flexible molecular species and other atomic groupings more generally. In particular the more complex sub-components of flexible molecules such as atomic bonds or angle based interactions etc.
Author contributions
H. L. Devereux: Methodology; Software; Validation; Formal analysis; Investigation; Data Curation; Writing - Original Draft; Writing - Review & Editing; Visualization. C. Cockrell: Methodology; Software; Validation; Formal Analysis; Investigation; Writing - Original Draft; Writing - Review & Editing. A. M. Elena: Conceptualization; Methodology; Software; Validation; Formal Analysis; Investigation; Resources; Data Curation; Writing - Original Draft; Writing - Review & Editing; Supervision; Project administration; Funding Acquisition. Ian Bush: Software; Validation; Data Curation; Writing - Review & Editing; Aidan B. G. Chalk: Software; Validation; Data Curation. Jim Madge: Software; Validation; Data Curation. Ivan Scivetti: Software; Validation; Data Curation. J. S. Wilkins: Software; Validation; Data Curation; Writing - Original Draft; Writing - Review & Editing. I. T. Todorov: Conceptualization; Methodology; Software; Validation; Formal Analysis; Investigation; Resources; Data Curation; Writing - Original Draft; Writing - Review & Editing; Supervision; Project administration; Funding Acquisition. W. Smith: Software; Conceptualization; Methodology; Validation; Data Curation. K. Trachenko: Conceptualization; Methodology; Validation; Formal analysis; Investigation; Resources; Writing; Writing-Review & Editing; Supervision; Project administration; Funding acquisition.
Data statement
The DLPOLY 5 source code is hosted on GitLab [11] (GPLv3.0). The input files and submission scripts to reproduce the statistical data for our on-the-fly methods can be found on Gitlab https://gitlab.com/apw951/dl_poly_workflows/-/tree/dlpoly_5_preprint?ref_type=tags.
Acknowledgements
We are grateful to EPSRC (for grant No. EP/W029006/1). This research utilised Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT http://doi.org/10.5281/zenodo.438045, STFC Scientific Computing Department’s SCARF cluster and the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick and funded by EPSRC GrantS EP/T022108/1 and the HPC Midlands+ consortium. Via our membership of the UK’s HEC Materials Chemistry Consortium, which is funded by EPSRC (EP/R029431), this work used the ARCHER UK National Supercomputing Service (http://www.archer.ac.uk)
Appendix A Kernels and derivation for SPME
Below follows a derivation for the generic kernels for order potentials.
As described in Essmann et al. [21] the general form of the function for is given as:
For the general case () we find:
where:
So for our simpler case:
Which, using a series expansion gives for even powers:
and for odd powers:
where:
which gives us the final form of the equation:
There are two simpler to solve cases, these are:
where :
and where is odd and
The general case of is harder to evaluate. Consider
This can be written
Which can easily be integrated by parts to give
This can be rearranged to give us an upwards recurrence
From which given I(0,x) and I(1,x) we can calculate all values of the integral for positive t. These starting integrals are straightfoward:
For negative t we need a downwards recurrence. This can also be obtained by rearranging the above
While we can use the above for even negative t it is not suitable for odd as stepping from to introduces a division by zero. We can solve this by noting
where is the first order exponential integral and using this as the starting point for odd negative .
From is trivially calculated
N.B. our is actually:
so we find that:
which gives:
as our actual stress kernel.
Appendix B Per-particle contributions for SPME
This section follows the notation of the original SPME paper [21]. , , label the lattice vectors. Note all is in the basis of the lattice vectors, should cartesian values be required a basis transformation is needed at the end (see the end of the force calculation for an example in the code)
Consider
where , , are the derivatives in , , .
It will be shown that all of the energy, force and stresses can be expressed in this form
Given the definition of the structure factor we can write:
Now applying the standard SPME methods.
First start working in scaled coordinates, where the scaling is by the size of the FFT grid we intend to use. So for a grid of size 1*2*3
Now applying the SPME approximation:
Apply periodic boundary conditions:
However we can go one step further. Differentiating A times on both sides w.r.t. and rearranging gives
The properties of cardinal B splines means this is easy to evaluate as
and since the splines are evaluated by recursion we already have all the terms we need already.
Inserting this into the above gives
Now the sum over is almost a discrete Fourier transform. The problems are missing out zero and the infinite range.
We can make it one if
1.
We define f(0)=0 – which it will be for a neutral cell
2.
We only sum out to -1. This is OK if f(m) has decayed to negligible values by the edge of the grid. This should be the case if we have converged the calculation carefully – ultimately the decaying potential term will kill off the function provided we have chosen the grid size and Ewald parameter correctly
Thus we can define a per particle contribution as
Appendix C Redefinition of control files
Old Style
P-cresol 340 temperature 340 ensemble nve 1.0 pressure 0.000 steps 10200000 multiple 1 print 2000 stack 100 stats 2000 trajectory 200000 2000 0 timestep 0.0005 cutoff 10.0 delr 0.5 rvdw cutoffΨ 10.000E+00 ewald precisionΨ 1.0000E-06 capΨΨ 1.0000E+04 shake toleranceΨ 1.0000E-05 quaternion tolerance 1.0000E-05 Ψ job timeΨ 2.5000E+05 close timeΨ 1.0000E+02 finish
New Style
title P-cresol 340 io_file_config CONFIG io_file_field FIELD io_file_statis STATIS io_file_revive REVIVE io_file_revcon REVCON io_file_cor COR io_file_currents CURRENTS print_frequency 2000 steps stats_frequency 2000 steps stack_size 100 steps vdw_cutoff 10.0 ang padding 0.125 ang cutoff 10.0 ang coul_method spme spme_precision 1e-06 ensemble nve temperature 340.0 K pressure_hydrostatic 0.0 katm timestep 0.0005 ps time_run 10200000 steps time_equilibration 0 steps restart clean traj_calculate ON traj_start 200000 steps traj_interval 2000 steps traj_key pos initial_minimum_separation 0.0 ang equilibration_force_cap 10000.0 k_B.temp/ang shake_tolerance 1e-05 ang
Appendix D Implementation of general correlations
In order to support arbitrary correlations in a modular and extensible manner, we separate the code into two components. First implementing the multi-tau algorithm [50], and second supplying an interface to submit data for correlation at runtime.
For the first, Algorithm 1 details updating a correlator with newly observed data and . In this listing the correlator structure includes rolling shift arrays for the data to be stored in, accumulators, and correlation values, all index by the block index in the hierarchical data structure. Every steps the algorithm proceeds to the next block by averaging the accumulated values of and that have previously been submitted. This step results in the increasing granularity when if . Every steps the shift arrays will again update from the start. The values of and may be real or complex (in which case a conjugate must be taken). In order to read out the correlation values, the implicit time lags due to any coarse graining need to be unwrapped as shown in Algorithm 2.
During a simulation multiple correlators are spawned, identified with correlating particular observables. As the simulation progresses, whenever a correlation must be updated the get method for the observables it tracks are called and the data submitted with (the first block).
For the second part, the interface calls for a pair of abstract Observable types to be supplied which each implement a get method. This method obtains the data to be correlated and acts as a kernel. By designing the module in this way a single loop over arbitrary pairs of Observable types may be completed to compute all correlations, without the need for additional selection logic. The specific functionality is implemented in types deriving from Observable as opposed to including this logic in the statistics collection subroutine itself. For some types this is a relatively trivial task. For example obtaining the velocity of a particular atom. Some a little more complex such as selecting one of the six current types which are also separated by atomic species and k-point. In all cases however, computation is separated from obtaining the data.